The body with mass m = 3 kg, under the action of the attractive force F = 10 N, slides in the horizontal plane and traverses the mug 20 m during the time 4 s, starting from the resting state. Calculate the friction force of the flesh exerted by the ground on the body.


1
Expert's answer
2020-12-14T07:17:28-0500

Find acceleration:


"a=\sqrt{2d/t^2}."

Describe the motion according to Newton's second law:


"ma=F-f,\\
f=F-ma=F-m\sqrt{2d/t^2}=5.3\text{ N}."